OpenStudy (anonymous):

Determine if the solution set for the system of equations shown is the empty set, contains one point or is infinite. x + y = 5 x + y = 7

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Lyre help me toooooooo

OpenStudy (lyrae):

Using normal substitution, or something from linear algebra (gauss-jordan elimination or determinant)?

OpenStudy (lyrae):

@feishdawwn

OpenStudy (lyrae):

Then I'll just do the most simple one: 1. x + y = 5 => y = 5 - x x + y = 7 x + 5 - x = 7 <=> 5 = 7 this is invalid and thus we have an empty set.